# Training Software Engineering Agents and Verifiers with SWE-Gym

### v1

@@ Abstract
We present SWE-Gym, the first environment for training real-world software engineering (SWE) agents. SWE-Gym contains 2,438 real-world Python task instances, each comprising a codebase with an executable runtime environment, unit tests, and a task specified in natural language. We use SWE-Gym to train language model based SWE agents, achieving up to 19% absolute gains in resolve rate on the popular SWE-Bench Verified and Lite test sets. We also experiment with inference-time scaling through verifiers trained on agent trajectories sampled from SWE-Gym. When combined with our fine-tuned SWE agents, we achieve 32.0% and 26.0% on SWE-Bench Verified and Lite, respectively, reflecting a new state-of-the-art for open-weight SWE agents. To facilitate further research, we publicly release SWE-Gym, models, and agent trajectories. [^1] https://github.com/SWE-Gym/SWE-Gym
@@ 1 Introduction
Figure 1: SWE-Gym enables scalable improvements for software engineering agents. Top: Training time scaling shows consistent performance improvements as we obtain more training trajectories, with no signs of saturation at 491 trajectories. We use temperature $t=0$. Bottom: For inference time scaling, we generate a number of candidate trajectories per task and select the best using a verifier trained on SWE-Gym. This approach demonstrates roughly logarithmic gains with the number of sampled solutions. $t=0$ (excluded from regression) is used as the first hypothesis to be consistent with the top figure; later rollouts use $t=0.5$.
@@
Table 1: SWE-Gym is the first publicly-available training environment combining real-world software engineering tasks from GitHub issues with pre-installed dependencies and executable test verification. Repository-level: whether each task is situated in a sophisticated repository; Executable Environment: whether each instance in the resource comes with an executable environment with all relevant dependencies pre-installed; Real task: whether the instruction for each instance is collected from human developers.
Language models (LMs) have remarkable promise in automating software engineering (SWE) tasks, as most clearly measured by recent progress on recent benchmarks like SWE-Bench (Jimenez et al., 2024) and Commit0 (Zhao et al., 2024). While LM-based SWE agents have shown significant performance gains through improving agent-computer interfaces (Yang et al., 2024) and prompting strategies (Wang et al., 2024c), advances in SWE agents have been limited by a reliance on proprietary models, rather than improving underlying LM itself.
Unlike other domains where supervised fine-tuning and reinforcement learning have significantly improved LM capabilities, for example in general chat (Ouyang et al., 2022), mathematical reasoning (Shao et al., 2024; Yuan et al., 2024), and web navigation (Pan et al., 2024), software engineering agents currently lack suitable training environments. Creating such an environment for SWE agents is uniquely challenging. Real-world software engineering requires interaction with an executable runtime that has been prepared with the appropriate software dependencies and reproducible test suites, among other requirements. These challenges are reflected in the existing resources (Tab. 1). For example, the SWE-Bench (Jimenez et al., 2024) training split contains only the solutions (git patches that solve the task), and lacks executable environments and reward signals. R2E (Jain et al., 2024) uses synthetic instructions that are very far from real-world problems, while datasets such as CodeContests (Li et al., 2022) and APPS (Hendrycks et al., 2021a) focus only on isolated tasks rather than realistic repository-level coding problems.
@@
To bridge this gap, we present SWE-Gym, the first training environment combining real-world software engineering tasks from GitHub issues with pre-installed dependencies and executable test verification. [^2] We discuss a concurrent work by Golubev et al. (2024) in §A. SWE-Gym contains 2,438 Python tasks sourced from 11 popular open-source repositories (Tab. 2), providing useful environments for training LMs as agents and verifiers.
@@
SWE-Gym supports training state-of-the-art open-weight SWE agents. Based on the OpenHands (Wang et al., 2024c) agent scaffold for general-purpose software development (§5.1), we fine-tune a 32B Qwen-2.5 coder model (Hui et al., 2024b) using only 491 agent-environment interaction trajectories sampled using SWE-Gym, and achieve substantial absolute improvements of +12.3% (to 15.3%) and +13.6% (to 20.6%) on SWE-Bench Lite and SWE-Bench Verified respectively (§3.2).
@@
SWE-Gym is effective across agent scaffolds. In another agent scaffold based on a SWE-Bench specialized workflow (Moatless, Örwall 2024, §5.1), we observe an improvement to 19.7% (32B model) and 10.0% (7B model) on SWE-Bench Lite through self-improvement, where the LM interacts with SWE-Gym, receives reward from it, and learns to improve itself through rejection sampling fine-tuning.
@@
SWE-Gym supports training verifier models that enable inference-time scaling. We use the test suite executions provided by SWE-Gym to determine whether sampled agent trajectories are successful or not. Given these samples, we train a verifier model (i.e., an outcome-supervised reward model Cobbe et al. (2021)) that estimates a trajectory’s probability of success, and thus enables inference-time scaling. Specifically, at inference time we sample multiple agent trajectories, and select the one with the highest estimated reward according to the verifier. This approach further improves the resolve rate to 32.0% (+11.4% absolute improvement) on SWE-Bench Verified (§4.1.1, Fig. 1 bottom) and 26.0% on SWE-Bench Lite (§4.1.2), establishing a new state-of-the-art among systems with publicly accessible weights (Tab. 9).
@@
Our baseline training and inference-time scaling methods on SWE-Gym yield continuously improved results with increasing compute (Fig. 1). In the training phase, performance scales with the number of sampled trajectories up to our current limit of 491 trajectories, suggesting that performance is currently limited by sampling compute rather than SWE-Gym’s size. Similarly, using the agent and verifier trained by SWE-Gym, the bottom panel shows that adding more compute during the inference time steadily improves the results.
@@ 2 SWE-Gym Environment
SWE-Gym comprises 2,438 real-world software engineering tasks sourced from pull requests in 11 popular Python repositories, with pre-configured executable environments and expert-validated test cases, constructed in close alignment with SWE-Bench (Jimenez et al., 2024). These repositories are separate from those used in SWE-Bench to avoid contamination. These tasks require SWE agents to develop test-passing solutions for real-world GitHub issues using provided codebases and executable environments. Such agents must map from natural language descriptions of the issue, as well as the initial state of the repository, to a pull request represented as a git patch. To solve the task, agents must edit or create files, and can execute code in the environment for debugging and testing.
The standard SWE-Gym includes 2,438 tasks. We also identify a subset of 230 tasks, SWE-Gym Lite, which contains generally easier and more self-contained tasks that are suitable for rapid prototyping, in alignment with SWE-Bench Lite (Jimenez et al., 2024).
To support future research in SWE agent development and automatic dataset synthesis, we also release a large set of Python GitHub issues without executable environments, which we refer to as SWE-Gym Raw. SWE-Gym Raw includes 66,894 instances spanning 358 Python repositories.
@@ 2.1 Dataset Construction
Identify Repositories To identify repositories as the basis of SWE-Gym, we first use SEART GitHub search [^3] https://seart-ghs.si.usi.ch/ to filter a list of initial repositories. Unlike SWE-Bench, which focuses on the top 5k most downloaded PyPI libraries (Jimenez et al., 2024), we select Python repositories that were created before 7/1/2022 and have more than 500 stars, with at least 300 lines of code, more than 500 pull requests (PRs) and 100 contributors. This results in 358 repositories.
Extract Training Instances from Repositories. We then use SWE-Bench’s instance extraction script to convert these repositories into SWE-Bench-style instances, each corresponding to a Github issue including the natural language description of the issue, a snapshot of the repository in which the issue was created, and a set of unit tests. Over the 358 repositories, we extract 64,689 task instances. We refer to this dataset as SWE-Gym Raw, which is over three times larger than the 19K instances gathered in previous work (Jimenez et al., 2024) and includes nearly ten times as many repositories.
While SWE-Gym Raw instances contain code, issue descriptions and the solution, they do not contain executable environments. And it’s unclear if the unit tests are effective in evaluating the correctness of a solution. We focus on a subset of repositories for which we semi-manually create such executable environments. Of the 358 total repositories collected, we select a subset of 11 repositories with high number of instances.
@@
Setup Executable Environments and Verify Instances An executable environment with pre-installed dependencies is essential for developing software engineering agents, as it most closely reflects an agent’s setup at deployment time, and enables incremental unit test feedback during development. However, configuring dependencies for a specific version of the codebase can be particularly challenging due to the absence of a universal installation method for Python packages. This issue is further compounded by backward compatibility problems where an older version of codebase requires an older version of a dependency, making environment setup even more difficult in old GitHub issues. Despite these challenges, ignoring these difficult environments might cause a distribution bias that decreases the utility of SWE-Gym, so we take on this challenge for the instances from these 11 selected repositories. We manually configure the dependencies for each task instance given on the relevant configuration files (e.g., requirements.txt), continuous integration scripts, or documentation provided in the repository snapshot at the time of the instance’s issue creation. We then run the execution-based validation script from SWE-Bench to verify that the gold patch (i.e., the code diff submitted by a human developer) passes more unit tests than the original code. We estimate that the entire process takes around 200 human annotation hours and 10k CPU core hours. After validation and filtering out failed instances, we obtain 2,438 unit-test-validated instances from 11 different repositories. We publicly release the pre-built Docker images for each instance to allow full reproducibility. Each image takes up an average of 2.6 GB, totaling 6 TB.
@@ 2.2 SWE-Gym Lite
Solving software engineering tasks is hard and computationally intensive, costing usually \$1 or more per task with frontier models (Wang et al., 2024c). To improve research efficiency via faster agent evaluation, Jimenez et al. (2024) introduce SWE-Bench Lite, a canonical subset of 300 instances from SWE-Bench. Following the SWE-Bench Lite filtering pipeline, [^4] For details on its construction process, see https://www.swebench.com/lite.html. we delineate the SWE-Gym Lite split, comprising 230 instances. Similar to SWE-Bench Lite, this subset excludes tasks that require editing more than one file, tasks with poorly described problem statements, those with excessively complex ground-truth code diffs, and tests focused on error message validation.
@@ 2.3 Dataset Statistics
Our analysis suggests that tasks and repositories included in SWE-Gym are more challenging than those included in SWE-Bench. Tab. 2 shows that overall, SWE-Gym has statistics similar to SWE-Bench, with several key differences. Codebases in SWE-Gym, on average, have relatively fewer files than SWE-Bench, but a similar number of lines of code. However, SWE-Gym has significantly more lines and files edited per gold patch when compared to SWE-Bench. Longer issue descriptions and a larger number of edits in the gold patch may loosely correlate with higher task complexity. Additionally, our experiments find consistently lower model performance on SWE-Gym compared to SWE-Bench.555§B.3 contains details of these experiments. Beyond models and scaffolds over-fitting to SWE-Bench, one explanation for this decreased performance on SWE-Gym could be our deliberate inclusion of sophisticated repositories, such as pandas and machine-learning libraries like MONAI.
3 Training LMs as SWE Agents with SWE-Gym
With SWE-Gym, we train language model agents. Our primary objective is to establish reasonable baselines for SWE-Gym using two agent scaffolds (OpenHands, Wang et al. 2024c, §3.2; Moatless Tools, Örwall 2024, §3.3) and validate the effectiveness of our dataset.
@@ 3.1 Setting
Agent Scaffolds. We experiment with two types of agents scaffolds –general-purpose prompting and specialized workflows, demonstrated through OpenHands CodeAct (Wang et al., 2024c) and MoatlessTools (Örwall, 2024).
Policy Improvement Algorithm. As a baseline, we employ a simple policy improvement algorithm: rejection sampling fine-tuning (a.k.a. filtered behavior cloning). This method is widely used as a baseline in language model and reinforcement learning literature (Zhou et al., 2024; Snell et al., 2022). At each iteration, a policy is sampled to interact with the environment. This policy may either be a stronger teacher policy or the student policy itself. Only trajectories that achieve rewards exceeding a predefined threshold are retained. The student model is then fine-tuned to mimic these high-reward trajectories using the standard negative log-likelihood loss. In our setting, the reward threshold is set to include only successful trajectories.
Evaluation Metrics. We perform evaluation on SWE-Bench Lite and SWE-Bench Verified (Jimenez et al., 2024). We measure: (1) Resolve Rate, RR, (%), the percentage of resolved problems; (2) Empty Patch, EP, (%), the percentage of trajectories where none of the code in the repository is edited. We use OpenHands’s remote runtime (Neubig & Wang, 2024) feature to parallelize evaluation (e.g., execute unit tests).
Training Setup. We use Qwen-2.5-Coder-Instruct (Hui et al., 2024a) 7B, 14B, and 32B variants as our base models. For hyper-parameters and details of the training runs, please refer to §B.2.

### v2

@@ Abstract
We present SWE-Gym, the first environment for training software engineering (SWE) agents. SWE-Gym contains 2,438 real-world task instances, each comprising a Python codebase with an executable runtime environment, unit tests, and a task specified in natural language. We use SWE-Gym to train language model based SWE agents, and achieve up to 19% absolute gains in resolution rate on the popular SWE-Bench Verified and Lite test sets. We also experiment with inference-time scaling through verifiers trained on agent trajectories sampled from SWE-Gym. When combined with our fine-tuned SWE agents, we achieve 32.0% and 26.0% on SWE-Bench Verified and Lite, respectively, reflecting a new state-of-the-art for open-weight SWE agents. To facilitate further research, we publicly release SWE-Gym, models, and agent trajectories.
@@ 1 Introduction
Figure 1: SWE-Gym enables scalable improvements for software engineering agents. Top: Scaling the amount of training data shows consistent performance improvements as we obtain more training trajectories, with no signs of saturation at 491 trajectories. We use temperature $t=0$ for evaluation. Bottom: For inference time scaling, we generate a number of candidate trajectories per task and select the best using a verifier trained on SWE-Gym. This approach demonstrates roughly log-linear gains with the number of sampled solutions. $t=0$ (excluded from regression) is used as the first hypothesis to be consistent with the top figure; later rollouts use $t=0.5$.
Language models (LMs) have remarkable promise in automating software engineering (SWE) tasks, as most clearly measured by recent progress on benchmarks like SWE-Bench (Jimenez et al., 2024) and Commit0 (Zhao et al., 2024). While LM-based SWE agents have shown significant performance gains through improving agent-computer interfaces (Yang et al., 2024) and prompting strategies (Wang et al., 2024c), advances in SWE agents have been limited by a reliance on proprietary models, with limited research to improve the underlying LM itself.
Unlike other domains where supervised fine-tuning and reinforcement learning have significantly improved LM capabilities, such as chat (Ouyang et al., 2022), math reasoning (Shao et al., 2024; Yuan et al., 2024), and web navigation (Pan et al., 2024), software engineering currently lacks suitable training environments, and creating environments is uniquely challenging. Real-world software engineering requires interaction with an executable runtime that has been prepared with the appropriate software dependencies and reproducible test suites, among other requirements. These challenges are reflected in the existing resources (Tab. 1). For example, the SWE-Bench (Jimenez et al., 2024) training split contains only solutions (git patches that solve the task), missing the step-by-step actions taken by the developer to create each solution, and executable environments and reward signals. R2E (Jain et al., 2024) uses synthetic tasks that are very far from real-world problems, while datasets such as APPS (Hendrycks et al., 2021a) focus only on isolated tasks rather than realistic repository-level coding problems.
Table 1: SWE-Gym is the first publicly available training environment combining real-world SWE tasks from GitHub issues with pre-installed dependencies and executable test verification. Repository-level: whether each task is situated in a sophisticated repository; Executable Environment: whether each task instance comes with an executable environment with all relevant dependencies pre-installed; Real task: whether task instruction is collected from human developers.
@@
To bridge this gap, we present SWE-Gym, the first training environment combining real-world software engineering tasks from GitHub issues with pre-installed dependencies and executable test verification. SWE-Gym contains 2,438 Python tasks sourced from 11 popular open-source repositories (Tab. 2), providing useful environments for training LMs as agents and verifiers.
@@
SWE-Gym supports training state-of-the-art open-weight SWE agents. Based on the OpenHands (Wang et al., 2024c) agent scaffold for general-purpose software development (§2), we fine-tune a 32B Qwen-2.5 coder model (Hui et al., 2024b) using only 491 agent-environment interaction trajectories sampled using SWE-Gym, and achieve substantial absolute improvements of +12.3% (to 15.3%) and +13.6% (to 20.6%) in resolution rate on SWE-Bench Lite and SWE-Bench Verified respectively (§4.2).
@@
SWE-Gym is effective across agent scaffolds. In another agent scaffold based on a specialized workflow (MoatlessTools; Örwall 2024; §2), we experiment with self-improvement, where the LM interacts with SWE-Gym, receives reward from it, and learns to improve itself through rejection sampling fine-tuning. This self-improvement boosts performance up to 19.7% on SWE-Bench Lite.
@@
SWE-Gym supports training verifier models to enable inference-time scaling. We use test suites included in SWE-Gym to determine whether sampled agent trajectories are successful or not. Given these samples, we train a verifier model (i.e., an outcome-supervised reward model; Cobbe et al., 2021) that estimates a trajectory’s probability of success. This enables inference-time scaling, where we sample multiple agent trajectories and select the one with the highest estimated reward according to the verifier. This further improves the resolution rate to 32.0% (+11.4% absolute improvement) on SWE-Bench Verified (§5.1.1; Fig. 1 bottom) and 26.0% on SWE-Bench Lite (§5.1.2), establishing a new state-of-the-art among systems with publicly accessible weights (Tab. 9).
@@
Our baseline training and inference-time scaling methods on SWE-Gym yield continuously improved results with increasing compute (Fig. 1). In the training phase, performance scales with the number of sampled trajectories up to our current limit of 491 trajectories, suggesting that performance is currently limited by the compute budget for sampling rather than the number of tasks in SWE-Gym. Similarly, using the agent and verifier trained by SWE-Gym, the bottom panel shows that using more compute during inference time steadily improves the performance.
2 Related Work
Agents that solve GitHub issues.
We focus on software engineering agents designed to automatically resolve GitHub issues within the SWE-Bench framework (Jimenez et al., 2024). These agents take a GitHub issue and its associated code repository as input and generate a valid code modification (i.e., a git diff patch) to address the issue. The correctness of these modifications is verified using a human-written test suite. Existing agent designs are categorized by the extent of human priors integrated into their workflows: Specialized workflows (Xia et al., 2024; Örwall, 2024; Zhang et al., 2024b; Chen et al., 2024) involve human-defined stages (e.g., localization, code editing, patch re-ranking), where a LM is iteratively prompted for each stage to produce the final result. This approach reduces the task horizon and minimizes the need for long-term planning. However, specialized workflows require significant human engineering, may not generalize to novel issue types, and can fail if intermediate steps encounter problems. In contrast, general-purpose prompting ((Yang et al., 2024; Wang et al., 2024c)) rely on LM’s ability to plan over long horizons and generate actions based on a history of interactions without heavily pre-defined workflows. While more flexible, general approaches demand higher capabilities from the underlying LM and can be computationally expensive due to multiple interaction rounds. The most successful existing SWE agents are built on proprietary language models like GPT-4 or Claude and utilize specialized workflows to overcome these models’ limitations. This contrasts with other sequential decision-making domains (Silver et al., 2017; Akkaya et al., 2019), where learning-based approaches, such as reinforcement learning, drive success by enabling systems to learn from interactions and rewards to develop task competence. A key barrier in the SWE agent domain is the lack of appropriate training environments. Our experiments show that SWE-Gym can be used to build strong learning-based agents, accelerating research in this area.
Environments for training software agents.
There is no existing dataset suitable for training software engineering agents. SWE-Bench (Jimenez et al., 2024) is widely used for evaluating software engineering performance, but its training split lacks executable environments and success signals present in the evaluation split, making it useful only for imitation learning approaches. HumanEval (Chen et al., 2021) is designed for standalone code generation tasks, akin to coding competitions. Therefore, it falls short of addressing the complex challenges inherent in real-world, repository-level software engineering tasks, which involve thousands of files, millions of lines of code, and tasks such as bug fixing, feature development, and system optimization. Similarly, R2E (Jain et al., 2024) is a small evaluation dataset with 246 instances and, due to its synthetic nature, lacks the realism and complexity in real-world software engineering scenario. Our proposed SWE-Gym instead uses real-world GitHub issues as task, and associated executable unit tests for evaluation. This results in realistic and complex task formulations, aligning closely with real-world challenges.
Post-training: From chatbots and reasoners to agents.
Post-training, which fine-tunes pre-trained language models using supervised or reinforcement learning, significantly improves model performance across various domains. Techniques like RLHF (Ouyang et al., 2022) have become standard for adapting language models into chatbots, improving both performance and alignment (Qwen Team, 2024). In math reasoning, datasets such as MATH (Hendrycks et al., 2021b) and GSM-8K (Cobbe et al., 2021) facilitate the training and evaluation of policy and verifier models (Cobbe et al., 2021; Wang et al., 2024a). Earlier works (Wang et al., 2024b; Chen et al., 2023; Zeng et al., 2023; Wu et al., 2024) demonstrate that distilling agent trajectories from stronger models improve weaker models. Recent studies (Xi et al., 2024; Zhai et al., 2024; Bai et al., 2024) explore self-improving methods, showing that reinforcement learning or rejection sampling fine-tuning guided by reward enables LMs to enhance themselves without more capable teachers.
However, post-training typically depends on expert demonstration data or training environments with reliable reward signals, which are largely absent in the software engineering domain. This has led to a reliance on prompting-based methods with proprietary language models. Our work addresses this gap with SWE-Gym, a training environment based on real-world software engineering tasks that uses expert-written tests as reward signals. Our experiments demonstrate that SWE-Gym can build strong SWE agents without prompt engineering.
@@ 3 SWE-Gym Environment
SWE-Gym comprises 2,438 real-world software engineering tasks sourced from pull requests in 11 popular Python repositories, with pre-configured executable environments and expert-validated test cases, constructed in close alignment with SWE-Bench (Jimenez et al., 2024). These repositories are separate from those used in SWE-Bench to avoid contamination. These tasks require SWE agents to develop test-passing solutions for real-world GitHub issues using provided codebases and executable environments. Such agents must map from natural language descriptions of the issue, as well as the initial state of the repository, to a pull request represented as a git patch.
We also identify a subset of 230 tasks, SWE-Gym Lite, which contains generally easier and more self-contained tasks that are suitable for rapid prototyping, in alignment with SWE-Bench Lite (Jimenez et al., 2024). To support future research in SWE agent development and automatic dataset synthesis, we also release SWE-Gym Raw, a large set of Python GitHub issues without executable environments (64,689 instances spanning 358 Python repositories).
@@ 3.1 Dataset Construction
Identify Repositories. We first use SEART GitHub search [^1] https://seart-ghs.si.usi.ch/ to filter a list of initial repositories. Unlike SWE-Bench, which focuses on the top 5k most downloaded PyPI libraries (Jimenez et al., 2024), we select Python repositories that were created before July 1, 2022 and have more than 500 stars, with at least 300 lines of code, more than 500 pull requests (PRs) and 100 contributors. This results in 358 repositories.
Extracting Training Instances from Repositories. We use SWE-Bench’s instance extraction script to convert these repositories into task instances, each corresponding to a GitHub issue including the natural language description of the issue, a snapshot of the repository in which the issue was created, and a set of unit tests. Over the 358 repositories, we extract 64,689 task instances. We refer to this dataset as SWE-Gym Raw, which is over three times larger than the 19k instances gathered in previous work (Jimenez et al., 2024) and includes nearly ten times as many repositories.
While SWE-Gym Raw instances contain code, issue descriptions, and the solution, they do not contain executable environments or a guarantee that its unit tests are effective in evaluating the correctness of a solution. Thus, we focus on 11 repositories with numerous instances and semi-manually create executable environments for them.
@@
Setup Executable Environments and Verify Instances.
Creating executable environments with pre-installed dependencies is crucial for developing software engineering agents, as it mirrors deployment settings and allows for incremental unit test feedback. Configuring dependencies for specific codebase versions is challenging due to the lack of a universal Python package installation method and backward compatibility issues, especially for older GitHub issues. Ignoring these environments could introduce distribution bias, diminishing SWE-Gym’s utility. To address this, we manually configure dependencies for each task instance using relevant configuration files (e.g., requirements.txt), CI scripts, or documentation from the repository snapshot at the time of issue creation. We then use SWE-Bench’s execution-based validation script to ensure that the gold patch (the human-submitted code diff) passes more unit tests than the original code. This process required approximately 200 human annotation hours [^2] Annotations are done by a subset of the authors. and 10,000 CPU core hours. After validation and filtering out failed instances, we obtained 2,438 unit-test-validated instances from 11 repositories. For full reproducibility, we publicly release pre-built Docker images for each instance, totaling 6 TB.
@@ 3.2 SWE-Gym Lite
Solving software engineering tasks is computationally intensive, costing usually \$1 or more per task with frontier models (Wang et al., 2024c). To improve research efficiency via faster agent evaluation, Jimenez et al. (2024) introduce SWE-Bench Lite, a canonical subset of 300 instances from SWE-Bench. Following the SWE-Bench Lite filtering pipeline, [^3] For details on its construction process, see https://www.swebench.com/lite.html. we delineate the SWE-Gym Lite split, comprising 230 instances. Similar to SWE-Bench Lite, this subset excludes tasks that require editing more than one file, tasks with poorly described problem statements, those with excessively complex ground-truth code diffs, and tests focused on error message validation.
@@ 3.3 Dataset Statistics
Our analysis suggests that tasks in SWE-Gym are on average harder than those included in SWE-Bench. Tab. 2 shows that SWE-Gym has statistics similar to SWE-Bench, with several key differences. Codebases in SWE-Gym, on average, have relatively fewer files than SWE-Bench, but a similar number of total lines of code. However, gold patches in SWE-Gym have significantly more lines and files edited when compared to SWE-Bench’s gold patches. Additionally, we find models have consistently lower performance on SWE-Gym compared to SWE-Bench.444§B.4 contains details of these experiments. Beyond models and scaffolds overfitting to SWE-Bench, the decreased performance on SWE-Gym may also be due to our inclusion of sophisticated repositories like pandas and MONAI.
4 Training LMs as Agents with SWE-Gym
We experiment with training language model agents using SWE-Gym. We use two agent scaffolds (OpenHands, Wang et al. 2024c, §4.2; Moatless Tools, Örwall 2024, §4.3).
@@ 4.1 Setting
Agent Scaffolds. Recent LM-based SWE agents comprise a base language model, and a set of tools and prompts this base model has access to. This set of tools and prompting strategies is referred to as an agent scaffold, and recent work has developed numerous scaffolds for different purposes (refer to §2 for examples). We experiment with two types of agent scaffolds: one for general-purpose prompting (OpenHands CodeAct; Wang et al. 2024c) and one for specialized workflows (MoatlessTools; Örwall 2024), which allows us to measure the efficacy of SWE-Gym across diverse deployment settings.
Policy Improvement Algorithm. We use SWE-Gym to improve the underlying LM for a given SWE agent. As a baseline, we employ a simple policy improvement algorithm: rejection sampling fine-tuning (a.k.a. filtered behavior cloning), where we fine-tune the base LM on success trajectories sampled from SWE-Gym.
Evaluation Metrics. We use the standard SWE agent benchmarks SWE-Bench Lite and Verified (Jimenez et al., 2024) for evaluation. We report (1) resolution rate (%), the proportion of resolved task instances, and (2) Empty Patch (%), the proportion of trajectories where none of the code in the repository is edited. We use OpenHands remote runtime (Neubig & Wang, 2024) to parallelize evaluation (e.g., execute unit tests).
Technical Details. For base LMs, we use Qwen-2.5-Coder-Instruct  (Hui et al., 2024a) 7B, 14B, and 32B. §B.2 contains training run details.